How do I identify the solution for -3x-4y=2 and 8y=-6x-4?
Answer by Math_Boss(45)

-3x-4y=2 (1)
8y=-6x-4 (2)
Rearrange: 8y=-6x-4; -6x-8y=4;
-3x-4y=2 (1)
-6x-8y=4 (2)(those equations are the same; if you divide by 2)
Coinciding lines have infinitely many solutions.
Ans: infinitely many solutions